OpenStudy (anonymous):

Solve using Substitution: 7x+2y=-8 8y=4x

OpenStudy (anonymous):

@jojo7

OpenStudy (anonymous):

ok first change first equation to equal y

OpenStudy (anonymous):

so y= -7/2x-4

OpenStudy (anonymous):

now take y and place into 8y=4x

OpenStudy (anonymous):

8 -7/2x-4=4x

OpenStudy (anonymous):

continuing to solve yu get -28x-32=4x

OpenStudy (anonymous):

-32=32x

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

now take -1 and put into top equation

OpenStudy (anonymous):

7(-1)+2y=-8

OpenStudy (anonymous):

-7+2y=-8 2y=-1 y=-1/2

OpenStudy (anonymous):

Thanks again!

OpenStudy (anonymous):

was ur question mult. choice

OpenStudy (anonymous):

Nope. :( its sample work.

OpenStudy (anonymous):

oh ok well ur welcome :)